(Exercise 6.43 (Algorithmic)) The time in minutes for which a student uses a center of a major with a mean of 38 minutes. Assume a student arrives at the terminal just as another student is beginning to work on the terminal. a. What is the probability that the wait for the second student will be 15 minutes or less (to 4 decimals)? uter terminal at the computer center of a major university follows an exponential probablity distribution b. What is the probability that the wait for the second student will be between 15 and 45 minutes (to 4 decimals)? c. What is the probability that the second student will have to wait an hour or more (to 4 decimals)?
in an article about investment alternatives, Money magazine reported that drug stocks provide a potential for long-term growth, with over 54% of the adult population of the United States taking prescription drugs on a regular basis. For adults age 65 and older, 8S% take prescription drugs reg larty For adults age 18 to 64. 47% take prescription drugs regularly. The age 18-64 age group accounts for 87.8% of the adult population (Statistical Abstract of the United States, 2008). e a. What is the probability that a randomly selected adult is 65 or older? b. Given an adult takes prescription drugs regularly, what is the probability that the adult is 65 or older?
Answer #1

6.43)

a) P(X<15)=1-e-x/\beta =1-e-15/38 =0.3261

b)P(15<X<45)=e-15/38-e-45/38 =0.3679

c)P(X>60)=e-60/38 =0.2062

4.45)

a)

P(adult is 65 or older)=1-0.878=0.122

b)P(adult is 65 or older |takes drug regularly) =0.122*0.85/(0.122*0.85+0.878*0.47)=0.2008